### 06 Jan 2024 v1

# Large Language Models as Visual Cross-Domain Learners

Abstract
Recent advances achieved by deep learning models rely on the independent and identically distributed assumption, hindering their applications in real-world scenarios with domain shifts. To address the above issues, cross-domain learning aims at extracting domain-invariant knowledge to reduce the domain shift between training and testing data. However, in visual cross-domain learning, traditional methods concentrate solely on the image modality, neglecting the use of the text modality to alleviate the domain shift. In this work, we propose Large Language models as Visual cross-dOmain learners (LLaVO). LLaVO uses vision-language models to convert images into detailed textual descriptions. A large language model is then finetuned on textual descriptions of the source/target domain generated by a designed instruction template. Extensive experimental results on various cross-domain tasks under the domain generalization and unsupervised domain adaptation settings have demonstrated the effectiveness of the proposed method.
1 Introduction
Deep models have achieved promising performance in various applications of computer vision when testing images are sampled from the same distribution of training data [26, 14, 58, 42]. However, in real-world applications, domain shifts [87] remain a significant challenge due to the distribution discrepancy between training and testing domains. For instance, an autonomous driving car should be able to handle adverse weather which does not appear during the training stage [13, 50, 66], posing a substantial hurdle for model generalization. To alleviate the domain shift, cross-domain learning [82, 92, 85] aims at extracting domain-invariant knowledge between source and target domains.
(a) CLIP embeddings.
(b) LLM embeddings.
Figure 1: t-SNE visualizations of samples from the Art Painting domain (marked by “o”) and the Cartoon domain (marked by “x”) on the PACS dataset. Samples of the same category are depicted in the same color. Implementation details are provided in Section 4.4.1.
Typical cross-domain learning scenarios include domain generalization (DG) [5, 75, 64] and unsupervised domain adaptation (UDA) [97, 85, 36, 95]. Specifically, DG trains a model from multiple source domains and evaluates it on an unseen target domain, while UDA assumes some unlabeled data from the target domain are available for training. Various DG [6, 40, 104] and UDA [85, 97, 45] methods have been proposed to alleviate the domain shift by learning domain-invariant features in the image modality. For example, under the DG setting, MIRO [6] proposes to learn features that are similar to the oracle representation. For the UDA setting, CDTrans [85] proposes a weight-sharing triple-branch transformer framework to enhance feature alignment and improve robustness to noisy pseudo-labels. However, the above methods only utilize image modality for cross-domain learning.
Instead of just using the image modality, recent efforts on large-scale vision-language models (VLMs) [55, 103, 21] have shown significant improvements in image classification by learning from massive paired image-text samples. For example, CLIP [55], which consists of an image encoder and a text encoder, is trained on 400 million image-text pairs by contrastive learning [8]. However, as shown in Figure 0(a), learning domain-invariant features in CLIP is challenging due to the large domain shift, leading to sub-optimal cross-domain performance. Large language models (LLMs) [70, 4] have impressive zero-shot capability for text-based tasks. Nevertheless, it is essential to acknowledge that the success of LLMs does not readily extend to pure vision and vision-language tasks, primarily due to the inherent disparities in modalities and task structures. To the best of our knowledge, there is no work that integrates LLMs with visual cross-domain learning.
In this paper, we propose LLaVO, using Large Language models as Visual cross-dOmain learners, to integrate the powerful LLMs into visual cross-domain learning. Specifically, we first use various VLMs (including CLIP [55] and BLIP [39]) to convert images into detailed textual descriptions (i.e., tags, attributes, and captions). As shown in Example 4.4.3 placed in the experimental section, these descriptions contain crucial image content information. As a byproduct, based on these descriptions, we release a series of datasets in text modality for text classification tasks and put the details in Appendix D. Next, we design an instruction template for querying the LLM to obtain images’ categories based on their descriptions. As shown in Figure 0(b), the domain shift of LLM embeddings is smaller than that of CLIP. Though LLMs perform well in natural language understanding, they may be distracted by irrelevant contexts [63, 88, 12] and the issue of hallucinations [74, 96, 67], leading to undesirable outputs. Especially in cross-domain learning scenarios, there are many domain-specific, but classification-irrelevant pieces of information [87]. To deal with these challenges, we finetune the LLM using training data comprised of question-answer pairs, where the proposed template creates questions, and the answers correspond to the class tokens of the respective images. In the UDA scenario, where the unlabeled target domain data are available, we use predictions of target domain data as pseudo-labels to finetune the LLMs for boosting performance. Moreover, LLaVO can use any off-the-shelf large-scale VLMs and LLMs. Extensive experiments are conducted on benchmark datasets under the DG and UDA settings to demonstrate that the proposed LLaVO method achieves state-of-the-art performance. Especially, compared with existing VLM-based methods, LLaVO performs better, showing the benefits of integrating LLMs into cross-domain learning.
Our contributions are summarized as follows.
  • We are the first to integrate LLMs with visual cross-domain learning. The proposed LLaVO uses VLMs as a bridge between image and text modalities and designs a question instruction template for the LLM to predict the category of an image based on its description.
  • The proposed LLaVO can alleviate the domain shift by finetuning the LLM on the designed question instruction template to improve the instruction following ability. In UDA, we finetune the LLM with pseudo-labels to further improve the cross-domain performance.
  • Extensive experiments on diverse DG and UDA tasks demonstrate that LLaVO consistently achieves state-of-the-art performance, surpassing existing VLM-based methods.
  • To facilitate further research, we release a series of datasets in the text modality, including PACS, VLCS, OfficeHome, and DomainNet.
2 Related Work
Figure 2: An illustration of the proposed LLaVO framework for both UDA and DG. For DG, the source images are converted into the text modality through the VLMs. Then, the extracted descriptions are used to finetune an LLM. For UDA, we use the target domain samples with pseudo-labels, along with the source domain samples with ground truth labels to finetune the LLM.
The goal of cross-domain learning [82, 92] is to learn a model from the source domain(s) with the possible help of the target domain to learn domain-invariant features and apply them to the target domain. Domain generalization and unsupervised domain adaptation are two typical scenarios in cross-domain learning.
Domain Generalization (DG). DG [5, 75, 6, 101] aims to improve the generalization capabilities of the models in new domains that are unseen during the training phase. DG is widely used in real-world applications [101, 15] and various methods have been proposed. For example, SWAD [5] and SAGM [75] find flatter minima to improve the generalization capabilities. Instead of learning domain-invariant features, MIRO [6] is proposed to learn features that are similar to the oracle representation.
Unsupervised Domain Adaptation (UDA). Besides a source domain with labeled data, in the UDA scenario [32, 44, 97], unlabeled data in the target domain are available during the training phase. CDTrans [85] proposes a weight-sharing triple-branch transformer framework to enhance feature alignment and improve robustness to noisy pseudo-labels. From a perspective of game theory, PMTrans [44] builds up an intermediate domain with an effective ViT [14]-based solution to smoothly bridge the source and target domains. Instead of explicitly aligning the source and target domains, MCC [32] introduces a general class confusion loss as a regularizer. ELS [97] adopts the domain adversarial training strategy, and mitigates the impact of label noise by encouraging the domain discriminator to output soft domain labels.
However, most existing DG and UDA methods focus solely on the image modality, leading to sub-optimal cross-domain effects as shown in our experiments. On the contrary, our LLaVO leverages both the image and text modalities.
VLMs and LLMs. VLMs have shown promising performance in learning visual representations [29, 49, 102, 1]. A notable example is CLIP [55], which is trained on 400 million image-text pairs and has been applied to various downstream tasks [21, 78, 94]. Recently, some works have applied CLIP to address domain shift problems [22, 9]. For example, DPL [94] learns a lightweight prompt generator to generate a domain-specific prompt for efficiently adapting CLIP to an unseen domain. CLIPood [64] further leverages the semantic relationships between classes in the text modality and proposes a novel model optimization strategy to improve the generalization performance. PADCLIP [36] proposes a catastrophic forgetting metric to adjust the learning rate and alleviate the catastrophic forgetting issue of CLIP. However, finetuning CLIP directly leads to a sacrifice in cross-domain generalization capability [64, 36].
On the other hand, LLMs show impressive generalization capability in various natural language processing tasks, e.g., topic classification [31, 4], mathematical reasoning [80, 99, 89, 30], and question answering [4, 79, 98]. Moreover, instruction-tuning methods [79, 10, 52] further improve the zero-shot and few-shot performance of LLM. Despite these advancements, LLMs cannot be directly applied to pure vision tasks and vision-language tasks due to the inherent disparities in modalities and task structure. To address this issue, some works have utilized VLMs as a bridge between the image and text modalities. For example, Menon et al. [49] leverage an LLM (i.e., GPT-3 [4]) to obtain descriptions for each category and employ CLIP to make predictions based on both the images and the extracted textual descriptions. LENS [2] designs a set of highly descriptive visual modules based on VLMs, enabling LLMs to tackle visual tasks. Different from those works that assume that all the data are in the same domain, the proposed LLaVO focuses on visual cross-domain learning.
3 Methodology
In this section, we present the proposed LLaVO method.
3.1 Overview of The Entire Framework
Suppose there is a set of $S$ source domains $\{\mathcal{D}_{i}\}_{i=1}^{S}$, where $\mathcal{D}_{i}=\{({\bf x}_{j}^{i},y_{j}^{i})\}_{j=1}^{n_{i}}$ and $n_{i}$ denotes the number of samples in $\mathcal{D}_{i}$. Let $\mathcal{D}=\cup_{i=1}^{S}\mathcal{D}_{i}$ denote the entire source domain training data. In cross-domain learning, we aim to learn a model from source domains and then generalize the model to a target domain $\mathcal{T}$ whose data distribution differs from that of the source domains. In DG, target domain data are unavailable during the training process, while in the training process of UDA, the unlabeled target domain data are available. Here we study the single-source UDA and multi-source UDA is left for future research, making $S$ equal to 1 in UDA.
To begin with, LLaVO generates textual descriptions of images using pre-trained VLMs. After obtaining textual descriptions, we finetune an LLM with the designed question instruction template that combines these descriptions and their category labels. Finetuning enables the LLM to follow the question instruction template and focus on the domain-invariant information related to the classification. The overall framework of LLaVO is shown in Figure 2 and the entire algorithm is shown in Algorithm 1 of Appendix C. In the following sections, we introduce those steps one by one.
3.2 Generating Textual Descriptions for Images
For an image ${\bf x}$, we use VLMs to convert it into textual descriptions. Previous works [49, 86] have demonstrated that textual descriptions can serve as accurate information for classification. Following [2], we use pretrained VLMs to extract textual descriptions from three aspects: (i) Tags; (ii) Attributes; and (iii) Captions.
To extract tags, a comprehensive tag vocabulary is constructed by integrating diverse sources, including multiple datasets such as [60, 38, 11, 53, 51, 3, 83, 33, 24, 41, 35, 34]. Specifically, the tags are wrapped by a standardized template “A photo of {tag}” and then fed to the text encoder of the CLIP model to obtain the embeddings. For each image ${\bf x}$, we compute the cosine similarity between its embedding extracted by the image encoder of the CLIP model and tags’ embeddings. The Top-$K$ similar tags are then used in the textual description.
However, tags are usually too brief to contain detailed descriptions of images. To deal with this issue, we use GPT-3 [4] to enrich the tags with attribute information. Specifically, for each tag, we prompt GPT-3 with the input “What are useful features for distinguishing a {tag} in a photo?” to generate the crucial attributes of visual characteristics for each tag. Similarly, all attributes are fed into the text encoder of the CLIP model to obtain the embeddings. For each image ${\bf x}$, we first calculate the cosine similarity between the embedding of ${\bf x}$, extracted by the image encoder of the CLIP model, and the embeddings of the attributes. Then the Top-$M$ attributes with the highest similarity scores are used in the textual description.
As tags and attributes are concept-level descriptions of images, they lack specific content of images. To this end, we further enrich the textual information by extracting captions of images. We use an image captioning model (e.g., BLIP [39]) to generate Top-$N$ captions for each image by stochastic sampling [16].
Based on the extracted tags, attributes, and captions as described above, we can obtain the textual descriptions as
$\mathbb{D}({\bf x})=$ “Tags: {Top-$K$ tags}
Attributes: {Top-$M$ attributes}
[…]
For illustration, Example 4.4.3 shows an example image and its corresponding extracted description.
3.3 Question Instruction
After extracting textual descriptions from images, we can use LLMs for classification. Specifically, we design a question instruction template for querying the LLM to its category. The template $\mathbb{T}({\bf x})$ is shown in Example 3.3, where $\{\mathbb{D}({\bf x})\}$ denotes the placeholder of the textual description of ${\bf x}$ and {Category Set} is the placeholder of the candidate categories. An example of the category set is shown in Appendix B.1.
Example 1: Template $\mathbb{T}({\bf x})$. Give the information about an image: $\mathbb{D}({\bf x})$}. According to the information, choose the most similar category from the given options: {Category Set}. ### Answer:
3.4 Finetuning
Zero-Shot LLM (ZS-LLM) [79] feeds $\mathbb{T}({\bf x})$ into LLMs and computes the sum of per-token log-likelihood for each category. The category with the highest log-likelihood is chosen as the predicted category. However, as shown in Figure 2(a), the performance of ZS-LLM is not satisfactory in classification. The limited performance could be due to two reasons: (1) The output of LLM is correlated to but still misaligned with the ground-truth category [100]. See Appendix A.2 for further discussion. (2) LLMs are susceptible to the influence of irrelevant context (e.g., domain-specific words) within the descriptions, which can negatively impact their classification accuracy [63].
To address the above issues and further exploit the capabilities of LLMs in classifying samples based on textual descriptions, we formulate the classification problem as a question-answer problem $(\mathbb{T}({\bf x}),y)$, where $y$, the class label of ${\bf x}$, is represented by a sequence of class tokens and denotes the true answer to the question $\mathbb{T}({\bf x})$.
3.4.1 Finetuning for DG Tasks
Under the setting of DG, inspired by [5, 19], we assume that there exists a global labeling function that generates labels for multiple domains. Thus we combine the training datasets of all the source domains into a training dataset $\mathcal{D}$. Then, instruction tuning is conducted on $\{(\mathbb{T}({\bf x}),y):({\bf x},y)\in\mathcal{D}\}$ with the next-token prediction objective. As LLMs have billions of parameters (e.g., LLaMA2 series have 7B, 13B, and 70B parameters), fully finetuning the LLM is computationally expensive. Therefore, we adopt the widely used parameter-efficient finetuning strategy (i.e., LoRA [28]). Let ${\boldsymbol{\theta}}$ denote parameters in the LoRA modules. Formally, the training objective of DG is to minimize the training loss as
$\mathcal{L}_{dg}({\boldsymbol{\theta}})=-\sum_{({\bf x},y)\in\mathcal{D}}\!\!\!\!\log p_{{\boldsymbol{\theta}}}\big{(}y|\mathbb{T}({\bf x})\big{)}.$ (1)
3.4.2 Finetuning for UDA Tasks
To make LLaVO more suitable for UDA tasks, we propose the utilization of LLaVO in conjunction with pseudo-labeling techniques. Pseudo-labeling strategies [65, 90] have enjoyed success in UDA by leveraging the unlabeled target domain data [36, 73]. Under the UDA setting, LLaVO combines target domain samples with pseudo-labels and source domain samples with ground truth labels to finetune the LLM. Similar to Eq. (1), the pseudo-labeling-based finetuning objective of UDA is to minimize the training loss as
$\!\!\!\!\mathcal{L}_{uda}({\boldsymbol{\theta}})\!=\!-\!\!\!\!\sum_{({\bf x},y)\in\mathcal{D}}\!\!\!\!\log p_{{\boldsymbol{\theta}}}\big{(}y|\mathbb{T}({\bf x})\big{)}\!-\!\!\sum_{{\bf x}\in\mathcal{T}}\log p_{{\boldsymbol{\theta}}}\big{(}\hat{y}|\mathbb{T}({\bf x})\big{)},\!\!$ (2)
where $\hat{y}$ denotes the pseduo-label for each target domain sample. Specifically, to obtain the pseudo-labels of target domain samples, the LLM is first finetuned using the source domain training data $\mathcal{D}$ with a loss function identical to Eq. (1). Subsequently, the fine-tuned LLM is employed to generate pseudo-labels for target domain samples.
3.5 Inference
After finetuning LLM with the designed question instruction template, the LLM can make predictions from the category set. Thus, we use the finetuned LLM to evaluate the samples in the target domain $\mathcal{T}$ for both DG and UDA settings. Specifically, given a testing sample ${\bf x}$, we feed the wrapped question $\mathbb{T}({\bf x})$ to the finetuned LLM to generate its answer as the prediction.
4 Experiments
In this section, we empirically evaluate the proposed LLaVO method.
4.1 Experiments on DG Tasks
Datasets. Experiments are conducted on four DG benchmark datasets: PACS [37], VLCS [17], OfficeHome [73], and DomainNet [17]. Due to page limit, the statistics of each dataset are offered in Appendix B.1. We follow the leave-one-out cross-evaluation protocol [23]: for each task, one domain is chosen as the target domain for testing, while the remaining domains serve as source domains for training.
Baselines. The proposed method is compared with existing DG methods including (1) ERM-ResNet [72], SWAD [5], MIRO [6], and SAGM [75] that use only the image modality. (2) Zero-shot CLIP (ZS-CLIP) [55], ERM-CLIP, DPL [94], and CLIPood [64] that use both the image and text modalities. ZS-CLIP uses the pre-trained CLIP to calculate the similarity between the query image embedding and the text embedding from the category set, and then choose the highest one as the prediction. For ERM-CLIP, following [81], we add a classification head after the image encoder and finetune both components in the source domain. Implementation details are provided in Appendix B.2.
Results. Tables 1, 2, 3, and 4 show the classification accuracies in the target domain for the PACS, VLCS, OfficeHome, and DomainNet datasets, respectively. As can be seen, LLaVO consistently achieves the best average accuracy on all the datasets. Moreover, LLaVO outperforms CLIP-based methods (i.e., ZS-CLIP, ERM-CLIP, DPL, and CLIPood), showing that using LLM could help VLMs in learning domain-invariant features to boost cross-domain performance. LLaVO also has good performance for each task on all the datasets. Specifically, LLaVO achieves the best performance across all tasks on the PACS, OfficeHome, and DomainNet datasets, and excels in three of four tasks on the VLCS dataset. Those results demonstrate the effectiveness of the proposed LLaVO method.
Table 1: Testing accuracies (%) of DG tasks on the PACS dataset. The best is in bold.
A C P S Avg
ERM-ResNet [72] 84.7 80.8 97.2 79.3 85.5
SWAD [5] 89.3 83.4 97.3 82.5 88.1
SAGM [75] 87.4 80.2 98.0 80.8 86.6
MIRO [6] 96.3 97.3 99.7 93.9 96.8
ZS-CLIP [55] 96.4 98.9 99.9 87.7 95.7
ERM-CLIP [72] 95.8 97.2 99.6 88.4 95.3
DPL [94] - - - - 97.3
CLIPood [64] 98.9 99.6 100.0 91.2 97.4
LLaVO 99.5 99.7 100.0 94.8 98.5
Table 2: Testing accuracies (%) of DG tasks on the VLCS dataset. The best is in bold.
C L S V Avg
ERM-ResNet [72] 97.7 64.3 73.4 74.6 77.5
SWAD [5] 99.8 63.3 75.3 79.2 79.1
SAGM [75] 99.0 65.2 75.1 80.7 80.0
MIRO [6] 98.4 64.5 80.3 83.7 81.7
ZS-CLIP [55] 99.7 61.8 70.1 73.9 76.4
ERM-CLIP [72] 97.0 68.9 82.8 85.1 83.5
DPL [94] - - - - 84.3
CLIPood [64] 99.7 68.1 79.9 90.8 84.6
LLaVO 99.9 69.9 81.3 88.1 84.8
Table 3: Testing accuracies (%) of DG tasks on the OfficeHome dataset. The best is in bold.
A C P R Avg
ERM-ResNet [72] 61.3 52.4 75.8 76.6 66.5
SWAD [5] 66.1 57.7 78.4 80.2 70.6
SAGM [75] 65.4 57.0 78.0 80.0 70.1
MIRO [6] 82.4 72.2 89.1 89.3 83.3
ZS-CLIP [55] 80.7 64.6 86.3 88.0 79.9
ERM-CLIP [72] 80.0 67.6 85.8 87.6 80.3
DPL [94] - - - - 84.2
CLIPood [64] 88.0 74.0 92.8 93.2 87.0
LLaVO 90.0 86.7 97.4 95.1 92.3
Table 4: Testing accuracies (%) of DG on the DomainNet dataset. The best is in bold.
C I P Q R S Avg
ERM-ResNet [72] 58.1 18.8 46.7 12.2 59.6 49.8 40.9
SWAD [5] 66.0 22.4 53.5 16.1 65.8 55.5 46.5
SAGM [75] 64.9 21.1 51.5 14.8 64.1 53.6 45.0
MIRO [6] 78.9 46.3 68.9 19.3 80.8 69.7 60.7
ZS-CLIP [55] 70.7 49.1 66.4 14.8 82.7 63.1 57.8
ERM-CLIP [72] 75.1 44.7 68.1 18.4 75.7 66.2 58.0
DPL [94] - - - - - - 56.7
CLIPood [64] 77.6 54.7 72.5 20.7 85.7 69.9 63.5
LLaVO 83.7 56.7 75.1 21.3 87.8 74.4 66.5
Table 5: Testing accuracies (%) of UDA tasks on the OfficeHome dataset. The best is in bold.
[…]
ERM [72] 44.1 67.1 74.3 53.3 62.0 64.5 51.9 38.9 72.9 64.5 43.8 75.4 59.4
DANN [20] 52.5 62.6 73.2 56.9 67.0 68.3 58.4 54.1 78.3 70.8 60.8 80.6 65.3
AFN [84] 52.6 72.4 77.0 64.9 71.1 72.9 64.1 51.3 77.8 72.2 57.5 82.1 68.0
CDAN [46] 54.2 72.2 78.3 62.0 71.4 72.4 63.0 55.7 80.7 74.7 61.2 83.7 69.1
SDAT [59] 58.2 77.5 81.4 66.1 76.5 76.4 63.7 56.7 82.5 76.0 62.1 85.2 71.9
MCC [32] 56.8 79.8 82.7 67.8 77.0 77.8 67.0 55.4 81.8 74.0 61.4 85.4 72.2
ELS [97] 57.8 77.7 81.6 66.6 76.7 76.4 62.7 56.7 82.1 75.6 62.9 85.4 71.8
CDTrans [85] 68.8 85.0 86.9 81.5 87.1 87.3 79.6 63.3 88.2 82.0 66.0 90.6 80.5
PMTrans [105] 81.3 92.9 92.8 88.4 93.4 93.2 87.9 80.4 93.0 89.0 80.9 94.8 89.0
PADCLIP [36] 76.4 90.6 90.8 86.7 92.3 92.0 86.0 74.5 91.5 86.9 79.1 93.1 86.7
LLaVO 85.4 96.6 94.1 90.3 97.1 94.4 87.9 85.7 94.5 90.1 85.5 97.3 91.6
4.2 Experiments on UDA Tasks
Datasets. For UDA, experiments are conducted on two benchmark datasets: OfficeHome [73] and DomainNet [17]. The statistics of those datasets are provided in Table 12 of Appendix B.1.
Baselines. The proposed method is compared with various UDA methods using the image modality only, including ERM [72], DANN [20], AFN [84], CDAN [46], SDAT [59], MCC [32], ELS [97], CDTrans [85], and PMTrans [105]. We also compare with PADCLIP [36] that uses both the image and text modalities. Implementation details are provided in Appendix B.2.
Results. Tables 5 and 6 show the testing accuracies of UDA tasks on the OfficeHome and DomainNet dataset, respectively. As can be seen, LLaVO consistently achieves superior average accuracy on all datasets. Compared with baselines based on the image modality only (i.e., ERM, DANN, AFN, CDAN, SDAT, MCC, ELS, CDTrans, and PMTrans), LLaVO performs better on average, demonstrating the effectiveness of using both the image and text modalities. LLaVO also has good performance for each task on all the datasets. Specifically, LLaVO achieves the best performance on 11 of 12 tasks and 24 of 36 tasks on the OfficeHome and DomainNet datasets, respectively. Moreover, LLaVO outperforms the CLIP-based method (i.e., PADCLIP), indicating that incorporating LLMs into cross-domain learners is more effective in cross-domain learning.
Table 6: Testing accuracies (%) of UDA tasks on the DomainNet dataset.
ERM C I P Q R S Avg MCC C I P Q R S Avg CDTrans C I P Q R S Avg
C - 18.0 36.2 12.1 53.6 42.6 32.5 C - 19.9 39.6 9.0 56.9 43.6 33.8 C - 29.4 57.2 26.0 72.6 58.1 48.7
I 41.0 - 35.7 4.7 52.9 31.2 33.1 I 37.2 - 38.1 3.0 54.8 26.6 31.9 I 57.0 - 54.4 12.8 69.5 48.4 48.4
P 46.0 18.5 - 6.2 59.9 38.5 34.4 P 48.4 19.7 - 4.4 61.1 41.2 35.0 P 62.9 27.4 - 15.8 72.1 53.9 46.4
Q 13.6 1.2 2.1 - 6.0 10.2 6.6 Q 18.5 3.9 9.2 - 17.6 13.0 12.4 Q 44.6 8.9 29.0 - 42.6 28.5 30.7
R 51.8 21.8 50.3 7.4 - 38.0 33.9 R 55.1 22.5 54.0 4.7 - 37.8 34.8 R 66.2 31.0 61.5 16.2 - 52.9 45.6
S 54.8 15.6 38.6 14.1 49.9 - 34.6 S 60.0 18.7 47.3 10.3 57.9 - 38.8 S 69.0 29.6 59.0 27.2 72.5 - 51.5
Avg 42.0 15.0 32.6 8.9 44.5 32.1 29.2 Avg 43.8 16.9 37.6 6.3 49.6 32.4 32.4 Avg 59.9 25.3 52.2 19.6 65.9 48.4 45.2
PMTrans C I P Q R S Avg PADCLIP C I P Q R S Avg LLaVO C I P Q R S Avg
C - 34.2 62.7 32.5 79.3 63.7 54.5 C - 55.1 71.1 36.8 84.2 68.1 63.1 C - 56.0 71.5 19.9 87.1 74.2 61.8
I 67.4 - 61.1 22.2 78.0 57.6 57.3 I 73.6 - 70.6 18.0 83.5 66.6 62.5 I 82.0 - 72.7 21.5 86.9 72.9 67.2
P 69.7 33.5 - 23.9 79.8 61.2 53.6 P 75.4 54.3 - 32.0 83.5 67.2 62.5 P 82.3 55.3 - 17.5 86.8 72.8 63.0
Q 54.6 17.4 38.9 - 49.5 41.0 40.3 Q 74.6 53.6 70.0 - 83.1 66.1 69.5 Q 79.2 52.5 66.3 - 84.8 70.7 70.7
R 74.1 35.3 70.0 25.4 - 61.1 53.2 R 76.4 54.9 72.7 31.7 - 67.5 60.6 R 83.7 55.7 74.2 18.8 - 73.6 61.2
S 73.8 33.0 62.6 30.9 77.5 - 55.6 S 76.3 54.9 71.7 34.9 83.6 - 64.3 S 83.1 55.3 75.4 21.4 87.1 - 64.4
Avg 67.9 30.7 59.1 27.0 72.8 56.9 62.9 Avg 75.3 54.6 71.2 30.7 83.6 67.1 63.7 Avg 82.1 55.0 72.0 19.8 84.2 72.8 64.7
Table 7: Effects of different textual information for DG tasks on the OfficeHome dataset. The best is in bold.
{NiceTabular}
cc—cccc c Tags+Attributes Captions A C P R Avg
✓ ✗ 85.5 81.9 96.8 94.0 89.6
✗ ✓ 88.2 83.3 95.3 92.8 89.9
✓ ✓ 90.5 85.8 97.4 95.4 92.3
Table 8: Effects of different LLM components for DG tasks on the OfficeHome dataset. The best is in bold.
A C P R Avg
Previous SOTA [64] 88.0 74.0 92.8 93.2 87.0
LLaVO (w/ Flan-T5-base) 85.0 81.7 92.8 90.4 87.5
LLaVO (w/ LLaMA2-7B) 90.0 86.7 97.4 95.1 92.3